5/22-23: MZ release from PSL
5/20-24: Exchanging cryo valves.
5/27-6/2: Y arm mass lock. (not done yet)
5/29: RT PC maintenance -> go to safe mode for all VIS connected Dolphin (specially BS)
6/3-6/8: Installation of TMS for ETMY
6/8: Engineering run
- Long term:
4/22-5/31: Yarm commissioning
- 5/27-6/30: DRMI commissioning
- 7/1-8/9: DRFPMI commissioning
- 8/12-10/4: DRFPMI ASC
- 10/7-11/15: Post Commissioning
- 11/18-12/13: Noise hunting
- Please reduce the number or speed of 16kHz channels.
Report of work
- COM: Tried PRMI, but BS yaw fluctuation was large about ~2urad.
- Type B: Increasing damping gain for IP yaw of BS did not work so much.
- DGS: Maintenance for exchanging RT-PC today, and fixing Dolphin.
- COM: PLL servo will be replaced during maintenance time.